OpenStudy (anonymous):

solve. show work: sqrt 8x+1=5 solve. show work: sqrt 8x+1=5 @Mathematics

OpenStudy (anonymous):

\[\Large \begin{array}{l} \sqrt {8x + 1} = 5\\ 8x + 1 = {5^2}\\ 8x = 24\\ x = \frac{{24}}{8} = 3 \end{array}\]

OpenStudy (anonymous):

x=3